OpenStudy (unklerhaukus):

@physics In an electromagnetic wave traveling through a vacuum, the Magnetic field H and Electric field E are said to be in phase, Is there any way to measure this effect?

OpenStudy (anonymous):

Do you mean observe this effect?

OpenStudy (unklerhaukus):

yeap

OpenStudy (unklerhaukus):

what is the difference?

OpenStudy (anonymous):

Measuring means you're looking for some device or construct to quantify it. Observe means that you merely want to see that it takes place.

OpenStudy (unklerhaukus):

well both would be nice

OpenStudy (anonymous):

I'm afraid I have no idea. Although I think light would be a great example since it travels through space.

OpenStudy (unklerhaukus):

But how do we know the E, H, fields are in phase,?

OpenStudy (anonymous):

Because, light propogates itself. Think about it, for something to move, a force is required. But once a photon starts moving, it has an electric field that generates a magnetic field that is perpendicular to the electric field. The resulting magnetic field, creates another electric field that is perpendicular to it and so on and so forth. It perpetually moves itself. This is assuming light is a wave of course. What are your thoughts?

OpenStudy (unklerhaukus):

That is a nice visual. However i still no reason that the fields are in phase. Wouldn't being in phase imply that the energy of the photon is fluctuating between a maximum energy and zero, What does this imply about conservation of energy ?

OpenStudy (anonymous):

Like I said, I don't know then. Any two waves would have to be atleast 0.5 wavelengths apart to fluctuate.

OpenStudy (unklerhaukus):

/? Sometimes E field and B field are zero, sometimes both are positive, sometimes both negative, Where does the energy go when both are zero?

OpenStudy (unklerhaukus):

Also in a standing electromagnetic wave in a vacuum the fields are π/2, 90° outta phase. I dont really understand why that is either

OpenStudy (anonymous):

Do you understand why it is from an analytical point of view? And just want a mechanism by which to measure it? Or are you just confused about why they would be in the first place?

OpenStudy (unklerhaukus):

Those reasons are the same thing

OpenStudy (anonymous):

I don't think so, but regardless... The time rate of change of the magnetic field is equal to the curl of the electric field. Have you ever taken the curl of a traveling wave \[ \vec{\nabla}\times <E_x,E_y,E_z>\sin(\vec{k}\cdot \vec{r} - \omega t)\] ? Try it out, and note the phase difference with the E field itself. This must be the time derivative of the B field, so..... it follows that E and B should be in phase.

OpenStudy (anonymous):

To answer your question as to where the energy goes, it's flowing! The energy density is not constant because the energy carried by the waves is propagating with them.

OpenStudy (anonymous):

Also @LordHades, that's not true, light doesn't propagate itself in the same way that a pulse down a rope that you've tied to a doorknob doesn't propagate itself. Electric fields and magnetic fields have no effect on photons.

OpenStudy (unklerhaukus):

I thought a photon Was a ElectroMagnetic wave. And im not sure how we can be certain the wave equations do not contain phase constants

OpenStudy (anonymous):

No. A photon is the massless particle that mediates the electromagnetic force. When you describe electricity and magnetism at a quantum mechanical level, you talk about photons. Classically, you talk about fields and electromagnetic waves. For many applications, the two treatments yield the same results, and since the classical picture is simpler, you learn it first. And phase constants don't just come out of nowhere for the same reason that \[\frac{d}{dx} \sin(x) \neq \cos(x+2) \]

OpenStudy (unklerhaukus):

\[{d \over dx }sin(x)=cos(x+2nπ) \]

OpenStudy (anonymous):

...which equates to no phase difference.

OpenStudy (unklerhaukus):

ok that does make sense. can you explain how the electric and magnetic field are in-phase the Standing electromagnetic wave

OpenStudy (anonymous):

Consider a cavity with conducting walls, so the transverse electric field is forced to be zero at the boundaries. The electric field now has the form \[\vec{E}(t) = \vec{E_0}\sin(\vec{k}\cdot \vec{r})\sin(\omega t) \] If you take the curl of this function the spatially varying part will change phase by pi/2 but the temporal part, the sin(w t), stays in phase now because it's not part of the same function. This means that the time derivative of the magnetic field is 90 degrees out of phase spatially and in phase temporally. This implies that the magnetic field itself is 90 degrees out of phase both spatially and temporally.